# Foliations by Critical Surfaces of the Hawking Energy in Asymptotically Flat Initial Data Sets

### v1

@@ Abstract.
Area-constrained critical surfaces for the Hawking quasi-local energy (“Hawking surfaces”) provide a natural setting for that energy: they enjoy positivity and rigidity properties, as shown in earlier work [27]. We construct large-scale foliations at infinity by Hawking surfaces in asymptotically Schwarzschild initial data sets. Using a Lyapunov–Schmidt reduction within the Willmore foliation framework of Eichmair–Koerber [9], we prove existence and uniqueness of the foliation and study its coordinate center. Under the dominant energy condition, we show that along the leaves of the foliation, the Hawking energy is positive and converges to the ADM energy in the large-sphere limit; moreover, subject to an explicit integral constraint, it is monotone along the foliation. Under weaker assumptions we construct an on–center family of Hawking surfaces that, while not necessarily a foliation, still enjoys positivity and the large–sphere limit. Finally, we obtain a rigidity statement and verify that our hypotheses hold in a broad class of data, initial data sets with harmonic or York asymptotics, thereby demonstrating the robustness of Hawking surfaces as a quasi-local energy tool in dynamical spacetimes.
@@ 1. Introduction and Main Results
Over the years many definitions of quasi-local energy have been proposed (see [31] for a comprehensive review), each with distinct strengths and limitations. To be physically acceptable, these definitions must satisfy certain physical conditions, particularly under the dominant energy condition, a realistic model for matter. These conditions include:
@@
    Small-sphere limit: On spheres shrinking to a point along a light cone, the quasi-local energy (after rescaling) should converge to the Einstein tensor evaluated on the timelike direction defining the cut. This reflects the fact that it captures the local energy density.
@@
Among quasi-local candidates, a prominent one is the one described by Hawking in [12], the so-called Hawking energy, which can be written for a surface $\Sigma\subset M$ as
@@
The Hawking energy is a natural and relatively simple quasi-local energy functional: it captures correct local behavior (e.g., the small-sphere limit, as shown by Horowitz and Schmidt [15], item (5) of the list above) but fails to be nonnegative on arbitrary surfaces. For instance, in Euclidean space every non-round 2-sphere has strictly negative Hawking energy, so one must single out a distinguished family of surfaces if one wants positivity.
The first landmark result in this direction is due to Christodoulou–Yau [5], who proved that, in the time-symmetric ($k=0$) setting under the dominant energy condition, the Hawking energy is nonnegative on constant mean curvature spheres. This suggests seeking a canonical class of surfaces, beyond the time-symmetric case, on which positivity (and ultimately rigidity, monotonicity, and the large-sphere limit) can be restored.
@@
Willmore surfaces have been extensively studied, and in the context of general relativity, a key development was made in [20] by Lamm, Metzger, and Schulze. They proved that an asymptotically Schwarzschild manifold admits a unique foliation near infinity by Willmore spheres—a “foliation at infinity” covering the end outside a compact set. In their framework, these surfaces are singled out as optimal for evaluating the Hawking energy since under nonnegative scalar curvature (dominant energy condition for $k=0$), the Hawking energy is nonnegative on Willmore surfaces, and it is monotonically nondecreasing along the foliation toward infinity.
Koerber [18] further showed that the leaves of the Willmore foliation are strict local area-preserving maximizers of the Hawking energy. Eichmair and Koerber later refined the construction of the foliation by Willmore spheres at infinity in [9], employing a Lyapunov-Schmidt reduction procedure. Moreover, in [8], they investigated the center of mass associated with this foliation.
@@
Previously studied in [11, 26, 27], these surfaces are well adapted to the Hawking energy: they have positivity and rigidity properties under the dominant energy condition [27] (items (1) and (2) of list above).
In the dynamical setting, local foliations by Hawking surfaces near a point were constructed in [26], where the small-sphere limit was also analyzed. At the other extreme, Friedrich’s Ph.D. thesis [10], building on the Willmore foliation framework [20], established large-scale foliations near infinity.
Beyond monotonicity and the large-sphere limit, an important complementary question for foliations at infinity is whether such a foliation also recovers a natural center of mass. In the time–symmetric case, Huisken–Yau showed that the CMC foliation yields the Hamiltonian center of mass [17], and Eichmair–Koerber proved the analogous statement for the Willmore foliation [8]. This prompts the question: under what additional symmetry or decay conditions does a foliation by Hawking surfaces become asymptotically centered, and does its coordinate center coincide, in the dynamical setting, with the spacetime (STCMC) center of mass introduced by Cederbaum and Sakovich [4]?
This paper has two goals: to extend the Eichmair–Koerber foliation [9] to dynamical data, and to establish the main physical properties of the Hawking energy along the resulting surfaces. We construct global foliations by Hawking surfaces in asymptotically Schwarzschild ends and analyze their principal features: existence, uniqueness, and asymptotic center. Along these foliations, the Hawking energy is positive, satisfies the large–sphere limit, admits a rigidity statement, and under further technical hypotheses is monotone. Under weaker assumptions, we also construct an on–center family of Hawking surfaces, which is not a foliation in general, but still satisfies positivity and the large–sphere limit. Taken together with the positivity and rigidity theory of [27] and the small–sphere limit of [15], our results show that, subject to technical assumptions, the Hawking energy satisfies the five main physical criteria for a quasi-local energy: four on Hawking surfaces and the small-sphere limit along light-cone cuts.
@@ 1.1. Main results
Notation. For functions $f,g$, we write
@@
if there are constants $C,\delta>0$ such that $|f(x)|\leq C\,|g(x)|$ for all $x\geq\delta$, and
@@
if $\lim_{x\to\infty}\frac{f(x)}{g(x)}=0$ (equivalently: for every $\varepsilon>0$ there exists $\delta>0$ with $|f(x)|\leq\varepsilon\,|g(x)|$ for all $x\geq\delta$). Also, $f(x)\leq\mathcal{O}(g(x))$ means that there exist $C,\delta>0$ such that $f(x)\leq C\,|g(x)|$ for all $x\geq\delta$.
@@
Let $(M,g,k)$ be a complete initial data set for the Einstein equations, where $M$ is diffeomorphic to $\mathbb{R}^{3}$ minus a compact set. We say $(M,g,k)$ is $\mathcal{C}^{4}$-asymptotically Schwarzschild of mass $m>0$ if there exists a compact set $K\subset M$ and a diffeomorphism
@@
We also obtain a uniqueness result for the previous foliation similar to the one obtained in [9].
Theorem 2.12. Under the same assumptions as the previous theorem. There exists a small constant $\epsilon_{0}>0$ and a compact set $K\subset M$ which depend only on $(M,g,k)$ such that the following holds. For every $\delta>0$, there exists a large constant $r_{0}>1$ such that every area-constrained critical sphere $\Sigma\subset M\setminus K$ of the Hawking functional with
@@
(7) $\mathrm{Sc}\geq-\mathcal{O}(\eta_{1}|x|^{-4})\quad\text{and}\quad\pi(\rho,\rho)=\mathcal{O}(\eta_{2}|x|^{-2})$
where $\pi=k-\operatorname{tr}k\,g$ and $\rho=\frac{x^{i}}{|x|}\partial_{i}$. Then there exists a compact set $K\subset M$, a number $\lambda_{0}>0$, and a family $\{\Sigma(\lambda):\lambda\in(0,\lambda_{0})\}$ of on-center stable area-constrained critical spheres $\Sigma(\lambda)$ of the Hawking functional, satisfying (4) with parameter $\lambda$ such that as $\lambda\to 0$, $|\Sigma(\lambda)|\to\infty$ and $R(\lambda)\to\infty$, where $R(\lambda):=\sup_{x\in\Sigma(\lambda)}|x|$.
@@
(8) $x^{i}\partial_{i}(|x|^{2}\mathrm{Sc})\leq\mathcal{O}(\eta_{1}|x|^{-2}),\quad k(\cdot,\rho)=\mathcal{O}(\eta_{2}|x|^{-2})\quad\text{and}\quad(\nabla_{\rho}\pi)(\rho,\rho)=\mathcal{O}(\eta_{2}|x|^{-3}),$
then the family of surfaces is unique in the sense of Theorem 2.12.
@@
We study the center of the foliations in a similar way to [8] for Willmore surfaces.
Theorem 3.7. Let $(M,g,k)$ be $\mathcal{C}^{4}$-asymptotic to Schwarzschild with mass $m>0$, satisfying the condition of the foliation Theorem 2.10. Furthermore, assume that
$|\mathrm{Sc}^{\text{odd}}|=\mathcal{O}(|x|^{-5}),\quad|k^{\text{even}}|+|x||(\nabla k)^{\text{odd}}|=\mathcal{O}(|x|^{-3})$
@@
Moreover, in Corollaries 3.9 and 3.11 we identify special cases where the foliation’s center converges more directly, and can be computed without reference to the intermediate spheres $S_{r}(r\xi(r))$ used in the construction. In particular, we find that the center of the foliation is more sensitive to the asymptotic distribution of the scalar curvature $\mathrm{Sc}$ and of the second fundamental form $k$ than the STCMC center of Cederbaum–Sakovich, and it need not coincide with the latter in general.
In Section 4 we prove the following monotonicity result.
@@
In conclusion, our analysis shows that, under the dominant energy condition, the Hawking energy on Hawking surfaces has the expected quasi-local features in a broad class of initial data sets. Monotonicity, however, currently holds only for foliations under an added integrability condition satisfied, for instance, in harmonic and York asymptotics. Moreover, as noted in [27], the rigidity results may rely on unnecessarily strong hypotheses, inducing an “excess positivity”. Relaxing these constraints could yield a sharper characterization.
Some of the material in Sections 2–4 first appeared in the author’s Ph.D. thesis [28].
@@ 2.1. Foliation construction
$W_{1}:=\Delta^{\Sigma}H+H|\mathring{B}|^{2}+H\,\mathrm{Ric}(\nu,\nu),\quad W_{2}:=P(\nabla_{\nu}\operatorname{tr}k-\nabla_{\nu}k(\nu,\nu))-2\,\operatorname{div}_{\Sigma}\!\big{(}P\,k(\cdot,\nu)\big{)}+\tfrac{1}{2}HP^{2},$
@@
For a normal variation $\frac{\partial F}{\partial t}=\alpha\nu$ one has the first-variation identity
(11) $\begin{split}\delta_{\alpha\nu}\left(\int_{\Sigma}P^{2}\,d\mu\right)&=\int_{\Sigma}2P\alpha(\nabla_{\nu}\operatorname{tr}k-\nabla_{\nu}k(\nu,\nu))+4Pk(\nabla^{\Sigma}\alpha,\nu)+\alpha HP^{2}d\mu\\ &=\int_{\Sigma}2W_{2}\alpha d\mu\end{split}$
If we now take a second normal variation $\frac{\partial\tilde{F}}{\partial s}=\tilde{\alpha}\nu$ and vary the right hand side of (11), we obtain the mixed second variation
@@
where $L_{J}\tilde{\alpha}:=-\Delta^{\Sigma}\tilde{\alpha}-\left(|B|^{2}+\mathrm{Ric}(\nu,\nu)\right)\tilde{\alpha}$ is the Jacobi operator, and we set $\nabla\pi(\nu,\nu)=(\nabla k)(\nu,\nu)-\nabla\operatorname{tr}k$ and $\nabla_{\nu}\nabla_{\nu}\pi(\nu,\nu)=(\nabla_{\nu}\nabla_{\nu}k)(\nu,\nu)-\nabla_{\nu}\nabla_{\nu}\operatorname{tr}k$. We denote this variation by
(13) $\delta_{\tilde{\alpha}\nu}\delta_{\alpha\nu}\left(\int_{\Sigma}P^{2}\,d\mu\right):=\int_{\Sigma}2DW_{2}(\alpha,\tilde{\alpha})\,d\mu,$
we will use this variation later.
We follow the notation of [9]. Given $\xi\in\mathbb{R}^{3}$ and $r>0$ we denote by $S_{\xi,r}$ the sphere
@@
Our foliation construction follows the strategy of Eichmair–Koerber, in particular their use of a Lyapunov–Schmidt reduction. It differs from the local foliations in [26] in that we project the equation only once, directly onto the complement of the kernel, rather than using an iterative projection scheme. Below, we give a brief outline of the construction.
$i)$ Perturbation of spheres. Begin with the large coordinate spheres $S_{r}(r\xi)$, $|\xi|<1$. An application of the Implicit Function Theorem produces a nearby surface $\Sigma_{r,\xi}$ of fixed area $4\pi r^{2}$ which satisfies the area-constrained Willmore equation up to the span of the first spherical harmonics.
ii)ii) Reduction to a finite-dimensional problem. It is shown that $\Sigma_{r,\xi}$ genuinely solves the constrained Willmore equation exactly if and only if $\xi$ is a critical point of the function
@@
In our setting we follow the same overall strategy, but now keep track of the extra terms involving the second fundamental form $k$. Moreover, we relaxed the conditions on the scalar curvature by changing the little $o$ bounds by a big $\mathcal{O}$ bound with a small constant $\eta_{i}$. A careful check of each step in [9] shows that their foliation continues to exist under these $\mathcal{O}$-type decays.
@@
with the $\mathcal{C}^{3,\alpha}$-topology. Let $\Lambda_{0}(S_{1}(0))$ and $\Lambda_{1}(S_{1}(0))$ be the constants and first spherical harmonics viewed as subspaces of $\mathcal{C}^{4,\alpha}(S_{1}(0))$, respectively. We use the symbol $\bot$ to denote the $L^{2}(S_{1}(0))$-orthogonal complements of these spaces. Now we derive the $k\neq 0$ generalization of the result [9, Lemma 16].
@@
where we denote $\hat{u}:=P_{(\Lambda_{0}(S_{1}(0))\times\Lambda_{1}(S_{1}(0)))^{\bot}}(u)$ the orthogonal projection of $u$ to $(\Lambda_{0}(S_{1}(0))\times\Lambda_{1}(S_{1}(0)))^{\bot}$, we also denote $\Lambda_{i}(S_{1}(0))$ by $\Lambda_{i}$ .
@@
Consider the map $\Theta_{\xi,r}(y)=r(\xi+y)$, like in [26] or [22], where local foliations in initial data sets were constructed, we consider rescaled data $g_{\xi,r}=r^{-2}\Theta^{*}_{\xi,r}g$ and $k_{\xi,r}=r^{-1}\Theta^{*}_{\xi,r}k$ on $M$. Then on this rescaled manifold, we define the function
@@
$|u_{\xi,r}|+r|^{\delta}\nabla u_{\xi,r}|+r^{2}|^{\delta}\nabla^{2}u_{\xi,r}|+r^{3}|^{\delta}\nabla^{3}u_{\xi,r}|+r^{4}|^{\delta}\nabla^{4}u_{\xi,r}|<c,\quad r^{3}|\lambda_{\xi,r}|<c$
@@
$|u|+r|^{\delta}\nabla u|+r^{2}|^{\delta}\nabla^{2}u|+r^{3}|^{\delta}\nabla^{3}u|+r^{4}|^{\delta}\nabla^{4}u|<\epsilon r,\quad r^{3}|\lambda|<\epsilon r,$
@@
We denote $u_{\xi,r}$ by $u$, $\lambda_{\xi,r}$ by $\lambda$, and $\Lambda_{\ell}(S_{\xi,r})$ by $\Lambda_{\ell}$ for $\ell=0,1$. Now just as in [9, Lemma 19] we obtain.
@@
There holds $P_{\Lambda_{0}}u=-2+\mathcal{O}(r^{-1})$ if $|\xi|<1-\delta$.
As in [9, Lemma 20] we expand $u$ in spherical harmonics.
@@
Where $P_{\ell}$ are Legendre Polynomials (see [9, Appendix B]).

### v2

@@ Abstract.
Area-constrained critical surfaces for the Hawking quasi-local energy (“Hawking surfaces”) provide a natural setting for that energy: they enjoy positivity and rigidity properties, as shown in earlier work [rigidiaz]. We construct large-scale foliations at infinity by Hawking surfaces in asymptotically Schwarzschild initial data sets. Using a Lyapunov–Schmidt reduction within the Willmore foliation framework of Eichmair–Koerber [eichko], we prove existence and uniqueness of the foliation and study its coordinate center. Under the dominant energy condition, we show that along the leaves of the foliation, the Hawking energy is positive and converges to the ADM energy in the large-sphere limit; moreover, subject to an explicit integral constraint, it is monotone along the foliation. Under weaker assumptions we construct an on–center family of Hawking surfaces that, while not necessarily a foliation, still enjoys positivity and the large–sphere limit. Finally, we obtain a rigidity statement and verify that our hypotheses hold in a broad class of data, initial data sets with harmonic or York asymptotics, thereby demonstrating the robustness of Hawking surfaces as a quasi-local energy tool in dynamical spacetimes.
@@ 1. Introduction and Main Results
Over the years many definitions of quasi-local energy have been proposed (see [Living] for a comprehensive review), each with distinct strengths and limitations. To be physically acceptable, these definitions must satisfy certain physical conditions, particularly under the dominant energy condition, a realistic model for matter. These conditions include:
@@
    Small-sphere limit: On spheres shrinking to a point along a light cone, the quasi‐local energy (after rescaling) should converge to the Einstein tensor evaluated on the timelike direction defining the cut. This reflects the fact that it captures the local energy density.
@@
Among quasi-local candidates, a prominent one is the one described by Hawking in [Hawma], the so-called Hawking energy, which can be written for a surface $\Sigma\subset M$ as
@@
The Hawking energy is a natural and relatively simple quasi-local energy functional: it captures correct local behavior (e.g., the small‐sphere limit, as shown by Horowitz and Schmidt [Haw], item (5) of the list above) but fails to be nonnegative on arbitrary surfaces. For instance, in Euclidean space every non-round 2-sphere has strictly negative Hawking energy, so one must single out a distinguished family of surfaces if one wants positivity.
The first landmark result in this direction is due to Christodoulou–Yau [Chriyau], who proved that, in the time‐symmetric ($k=0$) setting under the dominant energy condition, the Hawking energy is nonnegative on stable constant mean curvature spheres. This suggests seeking a canonical class of surfaces, beyond the time-symmetric case, on which positivity (and ultimately rigidity, monotonicity, and the large-sphere limit) can be restored.
@@
Area-constrained Willmore surfaces have been extensively studied, and in the context of general relativity, a key development was made in [willflat] by Lamm, Metzger, and Schulze. They proved that an asymptotically Schwarzschild manifold admits a unique foliation near infinity by Willmore spheres—a “foliation at infinity” covering the end outside a compact set. In their framework, these surfaces are singled out as optimal for evaluating the Hawking energy since under nonnegative scalar curvature (dominant energy condition for $k=0$), the Hawking energy is nonnegative on Willmore surfaces, and it is monotonically nondecreasing along the foliation toward infinity.
Koerber [Thomas] further showed that the leaves of the Willmore foliation are strict local area-preserving maximizers of the Hawking energy. Later, Eichmair and Koerber refined the construction of the foliation by Willmore spheres at infinity in [eichko], employing a different approach based on a Lyapunov–Schmidt reduction procedure, a method that we will also use in this paper. Moreover, in [willcen], they investigated the notion of center of mass naturally associated with this foliation.
In [willflat] Lamm–Metzger–Schulze construct the foliation through a direct perturbative analysis of large coordinate spheres in the Schwarzschild metric: the authors explicitly compute the first and second variations of the Willmore functional, derive a priori curvature and position estimates, and establish the invertibility of the linearized operator via comparison with the Schwarzschild case, allowing the use of the implicit function theorem to obtain existence and uniqueness under small perturbations of the metric. In contrast, Eichmair–Koerber [eichko] use a more abstract functional-analytic approach based on Lyapunov–Schmidt reduction, treating the area-constrained Willmore equation as a finite-dimensional critical point problem. By perturbing large spheres $S_{\lambda}(\lambda\xi)$ to nearby approximate solutions and identifying true solutions as critical points of a reduced functional $G_{\lambda}(\xi)$, they avoid the need for smallness assumptions, requiring only mild asymptotic conditions on the scalar curvature. This method yields broader existence and optimal uniqueness results, extending the foliation theory to a more general class of asymptotically Schwarzschild manifolds.
@@
Previously studied in [Alex, diaz2023local, rigidiaz], these surfaces are well adapted to the Hawking energy: they have positivity and rigidity properties under the dominant energy condition [rigidiaz] (items (1) and (2) of list above).
In the dynamical setting, local foliations by Hawking surfaces near a point were constructed in [diaz2023local], where the small-sphere limit was also analyzed. At the other extreme, Friedrich’s Ph.D. thesis [Friedrich2020], building on the Willmore foliation framework of Lamm, Metzger, and Schulze [willflat], established large-scale foliations near infinity using the same perturbative and implicit-function-theorem method developed in their work.
Beyond monotonicity and the large-sphere limit, an important complementary question for foliations at infinity is whether such a foliation also recovers a natural center of mass. In the time–symmetric case, Huisken–Yau showed that the CMC foliation yields the Hamiltonian center of mass [HY], and Eichmair–Koerber under more restrictive assumptions proved the analogous statement for the Willmore foliation [willcen]. This prompts the question: under what additional symmetry or decay conditions does a foliation by Hawking surfaces become asymptotically centered, and does its coordinate center coincide, in the dynamical setting, with the spacetime (STCMC) center of mass introduced by Cederbaum and Sakovich [STCMC]?
This paper has two goals: to extend the Eichmair–Koerber foliation [eichko] to dynamical data, and to establish the main physical properties of the Hawking energy along the resulting surfaces. We construct global foliations by Hawking surfaces in asymptotically Schwarzschild ends and analyze their principal features: existence, uniqueness, and asymptotic center. Along these foliations, the Hawking energy is positive, satisfies the large–sphere limit, admits a rigidity statement, and under further technical hypotheses is monotone. Under weaker assumptions, we also construct an on–center family of Hawking surfaces, which is not a foliation in general, but still satisfies positivity and the large–sphere limit. Taken together with the positivity and rigidity theory of [rigidiaz] and the small–sphere limit of [Haw], our results show that, subject to technical assumptions, the Hawking energy satisfies the five main physical criteria for a quasi-local energy: four on Hawking surfaces and the small-sphere limit along light-cone cuts.
@@ 1.1. Main results
Notation and asymptotic setting. We denote by $\delta$ and $\rho$ the Euclidean metric and radial direction $\frac{x}{|x|}$ respectively. For functions $f,g$, we write
@@
if there are constants $C,\hat{\delta}>0$ such that $|f(x)|\leq C\,|g(x)|$ for all $x\geq\hat{\delta}$, and
@@
if $\lim_{x\to\infty}\frac{f(x)}{g(x)}=0$ (equivalently: for every $\varepsilon>0$ there exists $\hat{\delta}>0$ with $|f(x)|\leq\varepsilon\,|g(x)|$ for all $x\geq\hat{\delta}$). Also, $f(x)\leq\mathcal{O}(g(x))$ means that there exist $C,\hat{\delta}>0$ such that $f(x)\leq C\,|g(x)|$ for all $x\geq\hat{\delta}$.
We denote the even and odd parts of $f$ by
and for tensors we take even/odd parts componentwise in the chosen chart. In what follows, all estimates are expressed in these coordinates, and all even/odd decompositions and $\mathcal{O}$–bounds are understood with respect to the asymptotically flat chart from the following definition.
@@
Let $(M,g,k)$ be a complete initial data set for the Einstein equations. We say $(M,g,k)$ is $\mathcal{C}^{4}$-asymptotically Schwarzschild of mass $m>0$ if there exists a compact set $K\subset M$ and a diffeomorphism
@@
(5) $x^{i}\partial_{i}(|x|^{2}\mathrm{Sc})\leq\mathcal{O}(\eta_{1}|x|^{-2}),\quad\mathrm{Sc}^{\text{odd}}=\mathcal{O}(\eta_{2}|x|^{-4})$
@@
We also obtain a uniqueness result for the previous foliation similar to the one obtained in [eichko].
Theorem 2.12. Under the same assumptions as the previous theorem. There exists a small constant $\epsilon_{0}>0$ and a compact set $K\subset M$ which depend only on $(M,g,k)$ such that the following holds. For every $\tilde{\delta}>0$, there exists a large constant $r_{0}>1$ such that every area-constrained critical sphere $\Sigma\subset M\setminus K$ of the Hawking functional with
@@
(7) $\mathrm{Sc}\geq-\mathcal{O}(\eta_{1}|x|^{-4}).$
Then there exists a compact set $K\subset M$, a number $\lambda_{0}>0$, and a family $\{\Sigma(\lambda):\lambda\in(0,\lambda_{0})\}$ of on-center stable area-constrained critical spheres $\Sigma(\lambda)$ of the Hawking functional, satisfying (4) with parameter $\lambda$ such that as $\lambda\to 0$, $|\Sigma(\lambda)|\to\infty$ and $R(\lambda)\to\infty$, where $R(\lambda):=\sup_{x\in\Sigma(\lambda)}|x|$.
@@
(8) $x^{i}\partial_{i}(|x|^{2}\mathrm{Sc})\leq\mathcal{O}(\eta_{1}|x|^{-2}),$
(9) $k(\cdot,\rho)=\mathcal{O}(\eta_{2}|x|^{-2}),\quad\pi(\rho,\rho)=\mathcal{O}(\eta_{2}|x|^{-2})\quad\text{and}\quad(\nabla_{\rho}\pi)(\rho,\rho)=\mathcal{O}(\eta_{2}|x|^{-3}),$
where $\pi=k-\operatorname{tr}k\,g$ and $\rho=\frac{x^{i}}{|x|}\partial_{i}$. Then the family of surfaces is unique in the sense of Theorem 2.12.
@@
In a similar way to [willcen] for Willmore foliations, we also study the center of the foliations constructed above. Because the foliation is composed of on-center spheres, any natural geometric center lies in the compact interior region $K$ from Definition 1.2, which is not visible through the asymptotic coordinate chart. For this reason, we introduce instead the coordinate center of the foliation, defined at infinity via the ambient chart.
Definition 1.4 (Coordinate center of a foliation).
$\vec{C}:=\lim_{\sigma\to\infty}\vec{c}(\Sigma_{\sigma}),$
Theorem 3.6. Let $(M,g,k)$ be $\mathcal{C}^{4}$-asymptotic to Schwarzschild with mass $m>0$, satisfying the conditions of the foliation Theorem 2.10. Furthermore, assume that
$\mathrm{Sc}^{\text{odd}}=\mathcal{O}(|x|^{-5}),\quad|k^{\text{even}}|+|x||(\nabla k)^{\text{odd}}|=\mathcal{O}(|x|^{-3})$
@@
Moreover, in Corollaries 3.8 and 3.10 we identify special cases where the foliation’s center converges more directly, and can be computed without reference to the intermediate spheres $S_{r}(r\xi(r))$ used in the construction. In particular, we find that the center of the foliation is more sensitive to the asymptotic distribution of the scalar curvature $\mathrm{Sc}$ and of the second fundamental form $k$ than the STCMC center of Cederbaum–Sakovich, and it need not coincide with the latter in general.
In Section 4 we study the nonnegativity and monotonicity of the Hawking energy along the foliation and prove the following result.
@@
In conclusion, our analysis shows that, under the dominant energy condition, the Hawking energy on Hawking surfaces has the expected quasi-local features in a broad class of initial data sets. Monotonicity, however, currently holds only for foliations under an added integrability condition satisfied, for instance, in harmonic and York asymptotics. Moreover, as noted in [rigidiaz], the rigidity results may rely on unnecessarily strong hypotheses, inducing an “excess positivity”. Relaxing these constraints could yield a sharper characterization.
Some of the material in Sections 2–4 first appeared in the author’s Ph.D. thesis [PenuelaThesis].
@@ 2.1. Foliation construction
(14) $W_{1}:=\Delta^{\Sigma}H+H|\mathring{B}|^{2}+H\,\mathrm{Ric}(\nu,\nu),\quad W_{2}:=P(\nabla_{\nu}\operatorname{tr}k-\nabla_{\nu}k(\nu,\nu))-2\,\operatorname{div}_{\Sigma}\!\big(P\,k(\cdot,\nu)\big)+\tfrac{1}{2}HP^{2},$
@@
For a normal variation $\frac{dF}{dt}=\alpha\nu$ one has the first-variation identity
(15) $\begin{split}\frac{d}{dt}\left(\int_{\Sigma}P^{2}\,d\mu\right)&=\int_{\Sigma}2P\alpha(\nabla_{\nu}\operatorname{tr}k-\nabla_{\nu}k(\nu,\nu))+4Pk(\nabla^{\Sigma}\alpha,\nu)+\alpha HP^{2}d\mu\\ &=\int_{\Sigma}2W_{2}\alpha d\mu\end{split}$
where we have used that $\frac{d}{dt}P=\alpha(\nabla_{\nu}\operatorname{tr}k-\nabla_{\nu}k(\nu,\nu))+2k(\nabla^{\Sigma}\alpha,\nu)$. If we now take a second normal variation $\frac{dF}{ds}=\tilde{\alpha}\nu$ and vary the right hand side of (15), we obtain, after grouping some terms, the mixed second variation
@@
where $L_{J}\tilde{\alpha}:=\frac{d}{ds}H=-\Delta^{\Sigma}\tilde{\alpha}-\left(|B|^{2}+\mathrm{Ric}(\nu,\nu)\right)\tilde{\alpha}$ is the Jacobi operator, and we set $\nabla\pi(\nu,\nu)=(\nabla k)(\nu,\nu)-\nabla\operatorname{tr}k$ and $\nabla_{\nu}\nabla_{\nu}\pi(\nu,\nu)=(\nabla_{\nu}\nabla_{\nu}k)(\nu,\nu)-\nabla_{\nu}\nabla_{\nu}\operatorname{tr}k$. We denote this variation by
(17) $\frac{d}{ds}\frac{d}{dt}\left(\int_{\Sigma}P^{2}\,d\mu\right):=\int_{\Sigma}2\mathcal{D}W_{2}(\alpha,\tilde{\alpha})\,d\mu,$
where we denote $\mathcal{D}W_{2}(\alpha,\tilde{\alpha}):=\frac{d}{ds}(W_{2}\alpha)+\alpha\tilde{\alpha}H^{2}P^{2}$.
We follow the notation of [eichko]. Given $\xi\in\mathbb{R}^{3}$ and $r>0$ we denote by $S_{\xi,r}$ the sphere
@@
Our foliation construction follows the strategy of Eichmair–Koerber [eichko], in particular their use of a Lyapunov–Schmidt reduction. It differs from the Lyapunov–Schmidt reduction of the local foliations in [diaz2023local] in that we project the equation only once, directly onto the complement of the kernel, rather than using an iterative projection scheme. Below, we give a brief outline of the construction in [eichko].
$i)$ Perturbation of spheres. Begin with the large coordinate spheres $S_{r}(r\xi)$, $|\xi|<1$. An application of the implicit function theorem produces a nearby surface $\Sigma_{r,\xi}$ of fixed area $4\pi r^{2}$ which satisfies the area‐constrained Willmore equation up to the span of the first spherical harmonics.
ii)ii) Reduction to a finite‐dimensional problem. It is shown that $\Sigma_{r,\xi}$ genuinely solves the constrained Willmore equation exactly if and only if $\xi$ is a critical point of the function
@@
In our setting we follow the same overall strategy, but now keep track of the extra terms involving the second fundamental form $k$. A careful check of each step in [eichko] shows that their foliation continues to exist if one relaxed the conditions on the scalar curvature by changing the little $o$ bounds by a big $\mathcal{O}$ bound with a small constant $\eta_{i}$. Therefore, in our results we consider such $\mathcal{O}$-type decays.
@@
with the $\mathcal{C}^{3,\alpha}$-topology. Let $\Lambda_{0}(S_{1}(0))$ and $\Lambda_{1}(S_{1}(0))$ be the constants and first spherical harmonics viewed as subspaces of $\mathcal{C}^{4,\alpha}(S_{1}(0))$, respectively. We use the symbol $\bot$ to denote the $L^{2}(S_{1}(0))$-orthogonal complements of these spaces. Now we derive the $k\neq 0$ generalization of the result [eichko, Lemma 16].
@@
The argument follows the same general scheme as in the time-symmetric case ($k=0$), but since it constitutes a central step in the analysis we present a complete version here for clarity.
@@
where we denote $\tilde{u}:=P_{(\Lambda_{0}(S_{1}(0))\times\Lambda_{1}(S_{1}(0)))^{\bot}}(u)$ the orthogonal projection of $u$ to $(\Lambda_{0}(S_{1}(0))\times\Lambda_{1}(S_{1}(0)))^{\bot}$, we also denote $\Lambda_{i}(S_{1}(0))$ by $\Lambda_{i}$ .
@@
Consider the map $\Theta_{\xi,r}(y)=r(\xi+y)$, like in [diaz2023local] or [main1local], where local foliations in initial data sets were constructed, we consider rescaled data $g_{\xi,r}=r^{-2}\Theta^{*}_{\xi,r}g$ and $k_{\xi,r}=r^{-1}\Theta^{*}_{\xi,r}k$ on $M$. Then on this rescaled manifold, we define the function
@@
$|u_{\xi,r}|+r|^{S_{\xi,r}}\nabla u_{\xi,r}|+r^{2}|^{S_{\xi,r}}\nabla^{2}u_{\xi,r}|+r^{3}|^{S_{\xi,r}}\nabla^{3}u_{\xi,r}|+r^{4}|^{S_{\xi,r}}\nabla^{4}u_{\xi,r}|<c,\quad r^{3}|\lambda_{\xi,r}|<c$
@@
$|u|+r|^{S_{\xi,r}}\nabla u|+r^{2}|^{S_{\xi,r}}\nabla^{2}u|+r^{3}|^{S_{\xi,r}}\nabla^{3}u|+r^{4}|^{S_{\xi,r}}\nabla^{4}u|<\epsilon r,\quad r^{3}|\lambda|<\epsilon r,$
@@
We denote $u_{\xi,r}$ by $u$, $\lambda_{\xi,r}$ by $\lambda$, and $\Lambda_{\ell}(S_{\xi,r})$ by $\Lambda_{\ell}$ for $\ell=0,1$. Now just as in [eichko, Lemma 19] we obtain.
@@
There holds $P_{\Lambda_{0}}u=-2+\mathcal{O}(r^{-1})$ if $|\xi|<1-\tilde{\delta}$.
As in [eichko, Lemma 20] we expand $u$ in spherical harmonics.
@@
Where $P_{\ell}$ are Legendre Polynomials (see [eichko, Appendix B]).